# Promptlinecolor=black,linewidth=2pt, leftmargin=1cm,rightmargin=1cm

### v2

@@ Abstract
In the era of big data, access to abundant data is crucial for driving research forward. However, such data is often inaccessible due to privacy concerns or high costs, particularly in healthcare domain. Generating synthetic (tabular) data can address this, but existing models typically require substantial amounts of data to train effectively, contradicting our objective to solve data scarcity. To address this challenge, we propose a novel framework to generate synthetic tabular data, powered by large language models (LLMs) that emulates the architecture of a Generative Adversarial Network (GAN). By incorporating data generation process as contextual information and utilizing LLM as the optimizer, our approach significantly enhance the quality of synthetic data generation in common scenarios with small sample sizes. Our experimental results on public and private datasets demonstrate that our model outperforms several state-of-art models regarding generating higher quality synthetic data for downstream tasks while keeping privacy of the real data,.
@@ 1 Introduction
Tabular data is the most common data format in high stake sectors like healthcare. There are many fundamental problems in dealing with tabular data, such as data scarcity, missing values, and irregularity. Among them, data scarcity problem has been the main roadblock. Many dataset in healthcare, such as clinical trial data, have small data size due to data collection cost and privacy risk, and consequently these data cannot afford modern machine learning (e.g., deep learning) which generally have thousands of parameters, at minimum.
@@
Recent advancements in generative models, particularly in text and image,(6; 23) have shown the benefits of technology for generating synthetic data that resembles the real data. Despite this potential, generating tabular data has not been fully tapped into; it has evolved through traditional statistical approaches, like Bayesian networks (29), to deep learning techniques, including autoencoders and Generative Adversarial Networks (GANs) (33). However, these methods requires large data for their training, contradicting our objective to solve data scarcity. Also, the sparsity and heterogeneity in tabular data makes GAN or other deep learning an suboptimal choice, as evidenced by the fact that the tree-based method (e.g., XGBoost) working better than deep learning model (9).
@@
Recently, advancements in large language models (LLMs) have also enabled researchers to use its general intelligence to synthesize tabular data.(5; 12) The premise is that prior knowledge encoded in the parameters of LLMs can provide contextual knowledge for coherent semantics that is required to learn the underlying data generation process. Several studies transformed tabular data to natural language via serialization, and used pre-trained LLMs to generate text containing the synthetic tabular data (5; 12; 17) . While fine-tuning LLMs has led to the creation of more nuanced synthetic data, this process again requires significant sample size, contradicting the objective of addressing data scarcity.
@@
In contrast, in-context learning presents a promising alternative by allowing LLMs to customize without compromising their general reasoning abilities (19) . Particularly, few-shot learning in in-context learning is to provide a few “examples” of data to let LLM to learn the patterns and mimic the examples (14). Our study aims to utilize this few-shot capability for synthetic tabular data generation. Currently, tabular generative model by in-context learning, such as (25) faces a critical shortcoming; it only accommodate too few “examples” (not an entire dataset available), thus discarding remaining data that cannot manage to fit into limited context length (input token). For example, in our real-world trial data ATACH2 with 37 variables, just ten samples consume 4,232 input tokens, whereas common input token size of LLMs, such as GPT3, is 2,048. This failure to utilize all available scarce data can make LLMs to perceive underlying data generating process merely based on “educated guess” with its prior knowledge, instead of data itself. This, consequently, cause distribution discrepancy between real data and synthetic data.
@@
Therefore, we aim to bridge this critical gap. Our key idea is make the data generation process explicit; the objective of our in-context learning is to generate better data generation process, as well as to generate individual data instances. Here the data generation process is a prompt text that consists of context of data and any simple “model” that describes relationship between data variables. We chose to use a Bayesian network or causal structure due to its interpretability and simplicity.
@@
However, another challenge is to identify ground-truth data generation process. Motivated by GAN’s adversarial training, we optimize the data generation process (“generator”) in adversarial training with “discriminator” (Table 1) . The discriminator’s role is to discriminates real data from the generated data, and we use the accuracy of the discriminator as a “loss function” to be minimized to optimize the generator. Unlike GAN, our generator is text format, which doesn’t have a derivatives. We address it by Optimization by Prompting, which leverages an independent LLM as an optimizer (34). After optimizing the data generation process, the LLM as generator use it to finally generate synthetic data.
@@
Optimizer Gradient descent Frozen LLM
@@
  • Few-shot synthetic data generation: Our model works when there are too few data to train a parametric model. It mitigates data scarcity problem in many small-size tabular data in healthcare.
@@
  • Explainability: Our LLM-based generator explicitly reveals data generation process and its reasoning, which are explainable by design. This enables transparency of our model and facilitate human feedback such as refining the knowledge.
@@ 2 Related Work
This section provides a brief overview of the most relevant prior work in the field of tabular data generation, LLM supported synthetic data generation, and different roles that LLM that play in real-world applications.
@@
Synthetic tabular data generation. Many studies have been proposed to generating high quality tabular data for privacy-preserving data sharing and augmenting training data size for machine learning models. Traditional method including Bayesian network (35; 29), approximation Bayesian computation (4), and SMOTE (7). Particularly, Bayesian network can represent a pairwise causal structure via directed acyclic graph (DAG) (21) , in which a directed edge between variable A and B exists if A causes B. Causal structure is a compact representation of underlying variable relationship, but does not fully capture the nonlinear numerical relationship between the different types of variables. Deep generative models have also been widely utilized. This include variational autoencoders (VAE) such as TVAE (33) and others (2; 31); GAN such as CTGAN (33), and diffusion model such as TabDDPM (15). The limitation of these model is that they requires sufficient sample size to train the generative model, which contradicts our aim to solve data scarcity.
@@
Synthetic data generation using LLMs. Recently, large language models emerged and has demonstrated its powerful performance in generating natural text. It has also shown a great potential in tabular data (8), such as for predicting (12; 10; 36; 17) and generating tabular data (5; 26; 37; 10). Among them, GReaT (5), the first model in this line, transformed the tabular data into text and fine-tuning LLM (GPT-2), including a feature order permutation step for added realism. However, these prior methods requires fine-tuning LLMs, which requires large data size and computing resources for its fine-tuning. This limitation of existing models motivates us to develop few-shot tabular data generative model.
@@
Multiple roles of LLM in real-world applications. In additional to typical natural text generation tasks it was originally trained for, LLMs have been utilized in various tasks. LLM has been used as an optimizer for data type that we can not calculate derivatives directly (34). This LLM as optimizer was used to optimize prompt (34) and optimize heuristic algorithms written in codes (24). LLMs also have been utilized for communicating with other LLMs (i.e., multi-agent LLM); multiple LLMs play different roles to accomplish a task collaboratively, such as coding,(11) question answering,(32) and online decision making.(27; 13) In this paper, we also leverage multiple LLMs with different roles to mimic adversarial training in GAN.
@@ 3.2.1 Generator
Note that subscript for iteration $i$ will be omitted for simplicity without loss of generalizability. Also, note that we used causal structure as a means to convey the relationship between variables within the prompt, thus obtaining ground-truth causal structure is not our ultimate goal.
@@
Conditional sampling. As MALLM-GAN generates synthetic data using LLM, it seamlessly inherits benefits of LLM such as conditional sampling. LLM predicts the next tokens given a user-provided context, even when the specified condition is rare or given as range. To conditionally sample the synthetic data, we modify the task instruction to contain specific conditions (Supplement Listing 4).
@@ 3.2.2 Discriminator
Based on the generated data, we evaluate and score the quality of $\theta$ by assessing how easy it is to distinguish generated synthetic data from real data. Naturally, this is a supervised learning task, rather than a reasoning task with LLMs. We use XGBoost to train $f$ because the tree-based ensemble models work best with tabular data. We build a discriminator $f$ such that $f:\mathcal{X}\to c$ where $\mathbf{x}\in\mathcal{X}$ and $f(\mathbf{x})$ is the predicted label $c$, which is 1 if $\mathbf{x}\in D_{train}$ and 0 if $\mathbf{x}\in D_{syn}$. Specifically, at each iteration $i$, a new set of synthetic data $D_{syn,i}$ is generated. We form the combined dataset $D_{train}\cup D_{syn,i}$. We assign labels to the combined dataset by
@@
We update the discriminator $f_{i}$ incrementally based on $f_{i-1}$. As the discriminator is a tree-based ensemble model, we incrementally add a new tree to the $f_{i-1}$ that fits the residuals of the current ensembles. We evaluate the accuracy of discriminator with $D_{test}$ and pass a pair of ($\theta_{i}$, $L(f_{i})$) to the optimizer where $L(f)$ denotes the discriminatory power of $f$ (e.g., accuracy, likelihood). We prefer to use accuracy (rather than likelihood) because this is a direct measurement that we aim to increase and our optimizer does not require numerical derivatives.
The discriminator obtain better discriminatory accuracy to distinguish real or synthetic data as the discriminator accumulate the discriminatory power of past iterations $10,...,i-$ 1 and also is updated with newly generated, more realistic synthetic data from current iteration $i$. However, on the other hands, as the $D_{syn}$ become more realistic over the iterations, it gets easier to fool the discriminator, and the discriminator’s accuracy decrease. Therefore, our discriminator obtain better discriminatory power during this adversarial optimization.
@@ 3.2.3 Optimizer
The next task is to optimize $\theta_{i}$ based on its score $L(f_{i})$. Our parameter to optimize is $\theta$, a text, which doesn’t have a derivatives. So we use Optimization by Prompting, which leverages LLM as an optimizer (34). To make LLM acts as a optimizer, we provide a meta-prompt, which consists of optimization task description such as “Your task is to optimize prompts for generating high-quality synthetic data. Aim … ” (see Example in Supplement listing 5 Line 3-6).
To leverage LLM’s in-context few-shot learning in the optimizer (34), we provide a few “examples” of possible solutions ($\theta,L(f)$). Note that the example here is different from data $(\mathbf{x},y)$. We keep the top $k$ solution pairs over the past iteration as the optimization trajectory to guide the optimization. We sort the score, so that the more desirable $\theta$ goes to the end of prompt. This will allows the LLM to recognize patterns among the data generation process with better score. See example in Supplement listing 5 Line 9-31.
@@ 6 D_syn = [LLM_generator(theta + example) for example in batch]
10 discriminator.fit(train, train_label)
@@ 4 Experiments
We present the evaluation results of MALLM-GAN. Our extensive experiments demonstrated that MALLM-GAN outperform baselines in generating high-quality synthetic data while preserving data privacy, thanks to the adversarial optimization of the data generation process. Additionally, MALLM-GAN’ provides explainable data generation through natural textual representation, effectively generating high-quality synthetic data based on user-provided conditions, even for rare categorical value or numeric ranges.
@@ 4.1 Setting
LLM. We used HIPPA-compliant Azure OpenAI GPT-3.5(6) as our generator and GPT-4 (20)(gpt-4-32k-0613) as our optimizer. Due to the extensive workload of the generator, we opted for the lighter, faster gpt-35-turbo-0125 model with a 16k context length. For the optimizer, requiring combinatorial search and high-level reasoning, we used the up-to-date gpt-4-32k-0613 model. As the optimizer requires more “creativity” compared to the generator, the generator’s temperature was set to 0.5, and the optimizer’s to 1.0 after multiple trials.
@@
Data. Our benchmarks include several datasets from various domains: two public datasets (Adult(3), Medical Insurance(1)), and two private medical datasets (ATACH2, ERICH) (18). To ensure fair comparison without memorization concerns (e.g., public datasets are in the training corpus of LLM), private datasets were included. Details can be found in Supplement Table 5.
Baselines. We compared MALLM-GAN against several state-of-art tabular generative models:
  • CTGAN (33): A GAN framework with a conditional generator for mixed data type.
  • TVAE (33): A variational autoencoder using the same preprocessing strategy as CTGAN, optimized by evidence lower-bound loss.
  • Be-GReaT (5): Transforms tabular data into text format and fine-tunes an auto-regressive generative LLM to model the distribution of real data.
@@
Other hyperparameters. Various techniques have been proposed to transform tabular data into natural language text (12). We tested several serializations, such as Manual Template (12) (“Age is 53, Work class type is self-employed, …”), which proved ineffective for moderate feature sizes; this serialization made the input prompt lengthy and talkative, only worked when feature size $|\mathbf{x}|$ is very small. Feature order permutation (5) also had negligible impact on performance. Similar to MALLM-GAN, prior work (25) uses in-context few-shot learning of pre-trained LLMs but incorporates post-hoc data selection, which is beyond our scope. A comparison without post-hoc selection is available in Table 3. Specific hyperparameters and computing resources are available in Supplement Section 2.
@@
Training data size vs. quality of synthetic data. We evaluated the impact of training data size $N=|D_{\text{train}}|$ on synthetic data quality by sampling subsets of different sizes ($N=100,200,400,800$). We particularly aimed to compare performances in low and moderate data size. For fair comparison between real and synthetic data, synthetic data was generated to match the size of real data ($|D_{train}|=|D_{syn}|$). We held out 100 samples as the test set $D_{test}$ before sampling and replicated experiments five times to estimate the standard error of the evaluation metrics. The batch size was set to be 40, with maximum iterations set to be 5, 4, 3, 2 for data sizes $N=100,200,400,800$, respectively.
4.2 Quantitative Evaluation

### v3

@@ Abstract
In the era of big data, access to abundant data is crucial for driving research forward. However, such data is often inaccessible due to privacy concerns or high costs, particularly in healthcare domain. Generating synthetic (tabular) data can address this, but existing models typically require substantial amounts of data to train effectively, contradicting our objective to solve data scarcity. To address this challenge, we propose a novel framework to generate synthetic tabular data, powered by large language models (LLMs) that emulates the architecture of a Generative Adversarial Network (GAN). By incorporating data generation process as contextual information and utilizing LLM as the optimizer, our approach significantly enhance the quality of synthetic data generation in common scenarios with small sample sizes. Our experimental results on public and private datasets demonstrate that our model outperforms several state-of-art models regarding generating higher quality synthetic data for downstream tasks while keeping privacy of the real data.
@@ 1 Introduction
Tabular data is the most common data format in high-stakes sectors like healthcare. There are many fundamental problems in dealing with tabular data, such as data scarcity, missing values, and irregularity. Among them, the data scarcity problem has been the main roadblock. Many datasets in healthcare, such as clinical trial data, have small data sizes due to data collection costs and privacy risks, and consequently, these data cannot afford modern machine learning (e.g., deep learning), which generally has thousands of parameters, at minimum.
@@
Recent advancements in generative models, particularly in text and image,(7; 28) have shown the benefits of technology for generating synthetic data that resembles real data. Despite this potential, generating tabular data has not been fully tapped into; it has evolved through traditional statistical approaches, like Bayesian networks (38), to deep learning techniques, including autoencoders and Generative Adversarial Networks (GANs) (42). However, these methods require large amounts of data for their training, contradicting our objective of solving data scarcity. Also, the sparsity and heterogeneity in tabular data make GAN or other deep learning a suboptimal choice, as evidenced by the fact that the tree-based method (e.g., XGBoost) works better than deep learning model (11).
@@
Recently, advancements in large language models (LLMs) have also enabled researchers to use their general intelligence to synthesize tabular data.(6; 14) The premise is that prior knowledge encoded in the parameters of LLMs can provide contextual knowledge for coherent semantics that is required to learn the underlying data generation process. Several studies transformed tabular data to natural language via serialization, and used pre-trained LLMs to generate text containing the synthetic tabular data (6; 14; 21). While fine-tuning LLMs has led to the creation of more nuanced synthetic data, this process again requires a significant sample size, contradicting the objective of addressing data scarcity.
@@
In contrast, in-context learning presents a promising alternative by allowing LLMs to customize without compromising their general reasoning abilities (24) . Particularly, few-shot learning in in-context learning is to provide a few “examples” of data to let LLM to learn the patterns and mimic the examples (17). Our study aims to utilize this few-shot capability for synthetic tabular data generation. Currently, tabular generative model by in-context learning, such as (31) faces a critical shortcoming; it only accommodate too few “examples” (not an entire dataset available), thus discarding remaining data that cannot manage to fit into limited context length (input token). For example, in our real-world trial data ATACH2 with 37 variables, just ten samples consume 4,232 input tokens, whereas the common input token size of LLMs, such as GPT3, is 2,048. This failure to utilize all available scarce data can make LLMs perceive the underlying data-generating process merely based on “educated guess” with its prior knowledge instead of the data itself. This, consequently, causes distribution discrepancy between real data and synthetic data.
@@
Therefore, we aim to bridge this critical gap. Our key idea is to make the data generation process explicit; the objective of our in-context learning is to generate a better data generation process, as well as to generate individual data instances. Here, the data generation process is a prompt text that consists of the context of data and any simple “model” that describes the relationship between data variables. We chose to use a Bayesian network or causal structure due to its interpretability and simplicity.
@@
However, another challenge is to identify the ground-truth data generation process. Motivated by GAN’s adversarial training, we optimize the data generation process (“generator”) in adversarial training with “discriminator” (Table 1). The discriminator’s role is to discriminate real data from the generated data, and we use the accuracy of the discriminator as a “loss function” to be minimized to optimize the generator. Unlike GAN, our generator is a text format, which doesn’t have derivatives. We address it by Optimization by Prompting, which leverages an independent LLM as an optimizer (43). After optimizing the data generation process, the LLM as generator uses it to finally generate synthetic data.
@@
Optimizer Gradient descent Frozen LLM and prompt
@@
  • Few-shot synthetic data generation: Our model works when there is too little data to train a parametric model. It mitigates data scarcity problem in many small-size tabular data in healthcare.
@@
  • Explainability: Our LLM-based generator explicitly reveals data generation process and its reasoning, which are explainable by design. This enables transparency of our model and facilitates human feedback, such as refining the knowledge.
@@ 2 Related Work
This section provides a brief overview of the most relevant prior work in the field of tabular data generation, LLM-supported synthetic data generation, and the different roles that LLM plays in real-world applications.
@@
Synthetic tabular data generation. Many studies have been proposed to generate high-quality tabular data for privacy-preserving data sharing and augment training data size for machine learning models. Traditional method including Bayesian network (44; 38), approximation Bayesian computation (5), and SMOTE (8). Particularly, a Bayesian network can represent a pairwise causal structure via directed acyclic graph (DAG) (26), in which a directed edge between variable A and B exists if A causes B. Causal structure is a compact representation of underlying variable relationship, but does not fully capture the nonlinear numerical relationship between the different types of variables. Deep generative models have also been widely utilized. This includes variational autoencoders such as TVAE (42) and others (2; 40); GAN such as CTGAN (42), and diffusion model such as TabDDPM (19). The limitation of these models is that they require a sufficient sample size to train the generative model, which contradicts our aim to solve data scarcity.
@@
Synthetic data generation using LLMs. Recently, large language models emerged and have demonstrated their powerful performance in generating natural text. It has also shown a great potential in tabular data (10), such as for predicting (14; 12; 45; 21) and generating tabular data (6; 32; 47; 12). Among them, GReaT (6), the first model in this line, transformed the tabular data into text and fine-tuning LLM (GPT-2), including a feature order permutation step for added realism. However, these prior methods require fine-tuning LLMs, which requires large data size and computing resources for fine-tuning. This limitation of existing models motivates us to develop few-shot tabular data generative model.
@@
Multiple roles of LLM in real-world applications. In addition to typical natural text generation tasks it was originally trained for, LLMs have been utilized in various tasks. LLM has been used as an optimizer for the data type that we can not calculate derivatives directly (43). This LLM as optimizer was used to optimize prompt (43) and optimize heuristic algorithms written in codes (29). LLMs also have been utilized for communicating with other LLMs (i.e., multi-agent LLM); multiple LLMs play different roles to accomplish a task collaboratively, such as coding,(13) question answering,(41) and online decision making.(35; 16)
LLM and causal structure discovery Causal structure discovery involves either data-driven or expert knowledge-driven approaches. Data-driven approaches discover causal structure from data by conditional independence tests (33; 34), score-based heuristics (36), or relaxing the discrete constraints into continuous optimization (48; 46). Despite these advances, identifying the ground-truth causal structure from data remains a significant challenge, particularly in complicated domains like healthcare or when data is scarce. Expert-driven approaches can be an alternative option, but these methods are time-consuming and require significant expert involvement. Recently, many studies (18; 23) suggest that LLMs, which encode prior knowledge in their parameters, can support causal discovery by complementing expert knowledge. In this paper, we leverage multiple LLMs with different roles to mimic adversarial training in GAN and use the heuristic causal structure discovery to guide the data generation process.
@@ 3.2.1 Generator
Note that subscription for iteration $i$ will be omitted for simplicity without loss of generalizability. Also, note that we used causal structure as a means to convey the relationship between variables within the prompt; thus, obtaining ground-truth causal structure is not our ultimate goal.
@@
Conditional sampling. As MALLM-GAN generates synthetic data using LLM, it seamlessly inherits the benefits of LLM, such as conditional sampling. LLM predicts the next tokens given a user-provided context, even when the specified condition is rare or given as a range. To conditionally sample the synthetic data, we modify the task instruction to contain specific conditions (Supplement Listing 4).
@@ 3.2.2 Discriminator
Based on the generated data, we evaluate and score the quality of $\theta$ by assessing how easy it is to distinguish generated synthetic data from real data. Naturally, this is a supervised learning rather than a reasoning task with LLMs. We build a discriminator $f$ such that $f:\mathcal{X}\to c$ where $\mathbf{x}\in\mathcal{X}$ and $f(\mathbf{x})$ is the predicted label $c$, which is 1 if $\mathbf{x}\in D_{train}$ and 0 if $\mathbf{x}\in D_{syn}$. Specifically, at each iteration $i$, a new set of synthetic data $D_{syn,i}$ is generated. We form the combined dataset $D_{train}\cup D_{syn,i}$. We assign labels to the combined dataset by
@@
We update the discriminator $f_{i}$ incrementally based on $f_{i-1}$. We evaluate the accuracy of the discriminator with $D_{test}$ and pass a pair of ($\theta_{i}$, $L(f_{i})$) to the optimizer where $L(f)$ denotes the discriminatory power of $f$ (e.g., accuracy, likelihood). We prefer to use accuracy (rather than likelihood) because this is a direct measurement we aim to increase and because our optimizer does not require numerical derivatives.
The discriminator obtains better discriminatory accuracy to distinguish real or synthetic data as the discriminator accumulates the discriminatory power of past iterations $10,...,i-$ 1 and is updated with newly generated, more realistic synthetic data from the current iteration $i$. However, on the other hand, as the $D_{syn}$ becomes more realistic over the iterations, it gets easier to fool the discriminator, and the discriminator’s accuracy decreases. Therefore, our discriminator obtains better discriminatory power during this adversarial optimization.
@@ 3.2.3 Optimizer
The next task is to optimize $\theta_{i}$ based on its score $L(f_{i})$. Our parameter to optimize is $\theta$, a text, which doesn’t have derivatives. So we use Optimization by Prompting, which leverages LLM as an optimizer (43). To make LLM acts as a optimizer, we provide a meta-prompt, which consists of the causal structure and the optimization task descriptions such as “Your task is to optimize prompts for generating high-quality synthetic data. Aim … ” (see Example in Supplement listing 5 Line 3-6).
To leverage LLM’s in-context few-shot learning in the optimizer (43), we provide a few “examples” of possible solutions ($\theta,L(f)$). Note that the example here is different from data $(\mathbf{x},y)$. We keep the top $k$ solution pairs over the past iteration as the optimization trajectory to guide the optimization. We sort the score, so that the more desirable $\theta$ goes to the end of prompt. This will allow the LLM to recognize patterns among the data generation process with better score. See example in Supplement listing 5 Line 9-31.
@@ 6 D_syn = [LLM_generator(theta + example) for example in batch]
10 discriminator.update(train, train_label)
@@
3.3 Comparison to GAN and convergence
MALLM-GAN’s adversarial training is motivated by GAN, but it differs fundamentally from traditional GANs in that it operates in a natural language optimization space, using LLMs for prompt-based generation, which lacks formal mathematical guarantees. Unlike gradient-based optimization in GANs, MALLM-GAN’s optimization relies on empirical refinement of prompts through adversarial optimization with a discriminator. Theoretical convergence analysis is challenging due to the absence of numerical gradients. However, empirical convergence, demonstrated in our experiments in Section 4 and prior work (43), shows stable convergence, where discriminator accuracy declines as prompts are refined. This practical convergence criterion serves as a reliable alternative to formal guarantees in real-world tasks.
@@ 4 Experiments
We present the evaluation results of MALLM-GAN. Our extensive experiments demonstrated that MALLM-GAN outperforms baselines in generating high-quality synthetic data while preserving data privacy, thanks to the adversarial optimization of the data generation process. Additionally, MALLM-GAN’ provides explainable data generation through natural textual representation, effectively generating high-quality synthetic data based on user-provided conditions, even for rare categorical values or numeric ranges.
@@ 4.1 Setting
LLM. We used HIPPA-compliant Azure OpenAI GPT-3.5(7) as our generator and GPT-4 (25)(gpt-4-32k-0613) as our optimizer. Due to the extensive workload of the generator, we opted for the lighter, faster gpt-35-turbo-0125 model with a 16k context length. For the optimizer, requiring combinatorial search and high-level reasoning, we used the up-to-date gpt-4-32k-0613 model. As the optimizer requires more “creativity” than the generator, the generator’s temperature was set to 0.5, and the optimizer’s to 1.0 after multiple trials.
Discriminator. Strong discriminators do not always contribute to a better generator (3). We tested Logistic regression, XGBoost, and neural network; we used the logistic regression model because it showed the highest performance while ensuring tractability during incremental updates over the iterations (Supplementary 6).
@@
Data. Our benchmarks include several datasets from various domains: three public datasets (Adult(4), Medical Insurance(1), Asia(30)), and two private medical datasets (ATACH2, ERICH) (22). To ensure fair comparison without memorization concerns of LLM (e.g., public datasets are in the training corpus of LLM), private datasets were included. Details are in Supplement Table 5.
Baselines. We compared MALLM-GAN with multiple state-of-the-art tabular generative models such as: traditional over-sampling techniques, SMOTE (9), the variational auto-encoder, TVAE (42), the generative adversarial network, CTGAN (42), LLM-based synthetic data generation model, Be-GReaT(6), and a diffusion model, TabDDPM (19). Similar to MALLM-GAN, a prior work (31) uses in-context few-shot learning of pre-trained LLMs but incorporates post-hoc data selection, which is beyond our scope. A comparison without post-hoc selection is available in Table 3.
@@
Other hyperparameters. Various serialization techniques have been proposed to transform tabular data into natural language text (14). We tested several serializations, such as Manual Template (14) (“Age is 53, Work class type is self-employed, …”), which proved ineffective for moderate feature sizes; this serialization made the input prompt lengthy and talkative, only worked when feature size $|\mathbf{x}|$ is very small. Feature order permutation (6) also had negligible impact on performance. Specific hyperparameters and computing resources are available in Supplement Section 2.
@@
Training data size vs. quality of synthetic data. We evaluated the impact of training data size $N=|D_{\text{train}}|$ on synthetic data quality by sampling subsets of different sizes ($N=100,200,400,800$). We particularly aimed to compare performances in low and moderate data size. For fair comparison between real and synthetic data, synthetic data was generated to match the size of real data ($|D_{train}|=|D_{syn}|$). We held out 200 samples as the test set $D_{test}$ before sampling and replicated experiments for each sub sample five times to estimate the standard error of the evaluation metrics. The batch size was set to be 50, with maximum iterations set to be 5, 4, 3, 2 for data sizes $N=100,200,400,800$, respectively.
4.2 Performance Evaluation